QUESTION 2 Find all values of z e C for which cos z-2/2 (10)

Thanks
Answer #1

z=a+ib\Rightarrow \cos(z)=\cos(a+ib)=\cos(a)\cos(ib)-\sin(a)\sin(ib)

Which is \cos(a)\cosh(b)-i\sin(a)\sinh(b)

Comparing we must have \cos(a)\cosh(b)-i\sin(a)\sinh(b)=-2\sqrt{2}i

And so \cos(a)\cosh(b)=0,\sin(a)\sinh(b)=2\sqrt{2}

We can never have \cosh(b)=0\Rightarrow \cos(a)=0

Thus \sin(a)=\pm 1

For \sin(a)=-1\Rightarrow \sinh(b)=-2\sqrt{2}\Rightarrow b=\sinh^{-1}(-2\sqrt{2})

And so a=2n\pi+\frac{3\pi}{2},b=\sinh^{-1}(-2\sqrt{2})

In case \sin(a)=1 we must have \sinh(b)=2\sqrt{2}\Rightarrow b=\sinh^{-1}(2\sqrt{2}) and so

a=2n\pi+\frac{\pi}{2},b=\sinh^{-1}(2\sqrt{2})

Thus, z=2n\pi+\frac{\pi}{2}+i\sinh^{-1}(2\sqrt{2}) OR z=2n\pi+\frac{3\pi}{2}+i\sinh^{-1}(-2\sqrt{2})

\blacksquare

Please do rate this answer positively if you found it helpful. Thanks and have a good day!